OpenStudy (anonymous):

Please Help!!! Find the derivative of f(x) = 12x^2 + 8x at x = 9. 256 -243 288 224

OpenStudy (anonymous):

\[f'(x)=24x+8\]\[f'(9)=24(9)+8\]\[f'(9)=224\]

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

Could you help me on one more?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

What happened to the x that went with the 8? And the other x from 24x^2?